OpenStudy (kkutie7):

HELP! An insulating spherical shell of inner radius R1 and outer radius R2 has a charge density \[ρ = ρ0r{2}/R^{2}-{1}\] a: Determine the electric field everywhere in space and graph |E~ | versus r. b: You have a conductor of radius R2. What surface charge density is needed for it to produce the same electric field for r > R2 as the insulating sphere does?

OpenStudy (kkutie7):



OpenStudy (kkutie7):

@satellite73

OpenStudy (michele_laino):

hint: since we have an electrostatic field, then we can apply this equation of Maxwell: \[\huge {\text{div}}{\mathbf{E}} = 4\pi \rho \left( r \right)\quad \left( {CGS} \right)\] please use the spherical polar coordinates, in order to make the needed computations

OpenStudy (kkutie7):

Never seen the equation before. I just thought that we have to go layer by layer. I was going to start with \[E_{r>R_{2}}\] and I read somewhere that it is like a point charge? That's as far as my understanding goes.

OpenStudy (anonymous):

Just ask for assistance here - submit your instructions to experts https://assignment.essayshark.com/blog/looking-for-physics-assignment-help/

OpenStudy (michele_laino):

What I wrote is a differential form of the Maxwell equation. Maybe you have seen such equation in integral form, namely: \[\huge \Phi \left( {\mathbf{E}} \right) = 4\pi {Q_{enclosed}}\] I have used the CGS system. Namely the flux of the electric field through a closed surface, is proportional to the electric charge inside or enclosed by such closed surface Please tag me and I will show the solution of your question, since I can not give direct answers @Kkutie7

OpenStudy (irishboy123):

this is all you need to know https://www.youtube.com/watch?v=QNIJC1emss8 http://www.maxwells-equations.com/gauss/law.php tag me if you need some more help

OpenStudy (irishboy123):

oh, and the total amount of charge in the hollow sphere can be gotten from a spherical integral: http://www.wolframalpha.com/input/?i=%5Cint_%7Bt+%3D+0%7D%5E%7B2+pi%7D+%5Cint_%7Bp+%3D+0%7D%5E%7B+pi%7D+%5Cint_%7Br+%3D+R_1%7D%5E%7BR_2%7D+(rho_o+%2F+R_1%5E2)+r%5E2+r%5E2+sin(p)+dr+dp+dt you will need make the upper limit a variable to get E inside the sphere.